# LLMs’ Reshaping of People, Processes, Products, and Society in Software Development: A Comprehensive Exploration with Early Adopters

### v1

@@ 3.1.3. Supporting Developers’ Personal Growth
  • Enhancing Reassurance, Confidence, and Independence: One theme that emerged in six interviews was LLMs’ role in providing developers with reassurance, which improved their confidence and independence. P2, a newer professional, shared his experience of comparing his answers to ChatGPT’s for corroboration: “It has made the [software engineering] process easier definitely. And the best part is I kind of get the reassurance of a lot of things — that if I’m doing it this way, I can ask it, and I can just check if ChatGPT is giving me a similar answer […], then I’m not doing anything which is majorly wrong. There will always be a few things here and there, but most of the parts would be correct. That’s what I get from it.
  • Improving Access to Information: A positive view of LLMs’ ability to ease access to information emerged in three interviews. P15 shared his belief that ChatGPT would democratize information access worldwide: “Now we have this tool that everybody in the world can use, that’s very, very inexpensive or free, that allows everybody, from all ages, from five years old to a hundred years old, to learn any topic that they want. I think a lot of countries are gonna be able to grow and thrive, knowing that they can now figure out how to grow these things, or what’s messed up with their economy, or analyzing these things […] You can’t say, ‘Oh, I grew up in a poor neighborhood. I couldn’t get the same information as the person who went to an Ivy League College,’ because that’s no longer the case. Now it comes down to whether or not you yourself have the ambition to learn the things that you say, that you would learn if education was free because now it’s free.
  • Enhancing Job Satisfaction: One participant, P14, highlighted ChatGPT’s role in his increased job satisfaction. As a non-native English speaker, he shared how ChatGPT helped him in areas he struggled with, thus allowing him to focus on what he cared about: “I feel like I’m very bad at writing essays in English. I feel like [ChatGPT] can write better essays than I can, and work my ideas better than I can in English. So it can help me in that way, that I don’t need to get my skills on some level. So I can write some okay essays or emails […] and I can focus more time on doing something that I enjoy and going deeper there, which it encourages me. And if I can focus more time [on] doing something that makes me happy and that I’m interested in, I can learn those things way faster and be way happier than if I have to learn something that I don’t really care about.
@@
  • Inability to Replace Human Decisions: Six participants noted LLMs’ inability to replace human interactions and decisions. For instance, P12 emphasized that software engineering involves more than just coding, highlighting the importance of human communication and collaboration: “Despite what [LLMs] can do, at the end of the day, it can’t replace — like, what’s difficult about being a software engineer isn’t coding particularly. Yeah, that’s part of it, [but] I think what’s difficult about being a software engineer, at least with what I do, is the communication that happens between teams, between coworkers, the email threads, and the chat threads that exist.
    Similarly, P5 stressed the ongoing necessity of human involvement in software development, especially in decision-making and understanding business requirements: “I still believe that you’ll still need a human person running the shots and doing the code. You still need someone to run ideas off of and to handle all of these human elements because you’d still need to take in the business case from a person and from the whole other team, and that’s a whole other topic and conversation. But there still is room in this world for the human interaction and human developer, in my perspective.
@@
  • Concerns about Dependency: Two participants mentioned their preference for solving problems on their own before asking LLMs. This was primarily done to avoid over-reliance—as P16 shared, “I try my best to at least spend some time on my own to figure it out because I don’t want to be too dependent on [ChatGPT].
  • Slower Implementation than Humans: Differences between the time it took humans and LLMs to create code were noted in one interview. P12 highlighted the efficiency of human problem-solving in cases where ChatGPT fails to understand or provide accurate solutions: “At some point, I’ll be like, ‘Okay, it just doesn’t understand me.’ And so I’ll give up and then just do it myself. Honestly, sometimes I feel like doing it myself can often be the easier answer — like, the actual, quickest path to getting what I want sometimes when it starts making these mistakes initially.
3.1.4. Assisting Developers’ Non-Technical Tasks
  • Consulting and Decision-Making: There were eleven participants discussing the use of LLMs for consultation or direction. For example, P7 highlighted their utility in uncertain situations: “If there is a question, [like] I’m not sure which tool I should use, I could probably just ask an open question to the ChatGPT like, ’Hey, could you give me some directions or some potential solutions to this situation?’ And ChatGPT could probably show me some high-level ways […] But with some traditional search engines, like Google, it’s kind of hard because if I don’t ask a specific question, or if I don’t ask for some specific tool, they cannot give me a suggestion or they give me some way unrelated suggestions.” In another case, P8 mentioned ChatGPT’s impact on resolving team debates: “[A]lways in technical teams, there exists debates on choosing options, options A and B — both are correct, but which one is better? In this way, we have a judge; we have someone that tells us which approach is better. In this way, I think it changed the whole software engineering [process] for me, that whenever we have a discussion with our colleagues in our team, in some cases, at least we have someone [ChatGPT] who says the final word.
  • Summarizing Text and Documentation: Nine participants mentioned the use of LLMs for summarizing different textual data, such as articles, papers, and documentation. With regard to reading documentation, P7 shared, “When I look into some documentations, instead of reading the whole thing — because I probably don’t need all of these things, I only need some piece of the information — I can ask ChatGPT, ‘Hey, read through this link and give me the summary.’
  • Supporting Internal Communication: The use of LLMs to facilitate internal communication by composing documents like quarterly updates, aiding in presentations by providing outlines and content suggestions, and assisting in explaining complex tasks to team members emerged in five interviews. For instance, P7 stated, “Since English is not my first language, [it] previously took me a lot of time to write that. But with ChatGPT, I can just show it this is what I’ve done. ‘Could you generate this kind of solution, this kind of document for me?’” P4 shared how he used ChatGPT to create presentations and demonstrations: “I also do a lot of presentations. So, [ChatGPT helps with] communicating what I’m building and what our team is working on internally to other teams, how to present that information, help me with slides, help me with what would make it a good demo, video, et cetera. It’s able to just give me outlines for that type of stuff!
@@ 3.1.4. Assisting Developers’ Non-Technical Tasks
  • Limited Summarizing and Explanation Capabilities: Two participants expressed dissatisfaction with LLMs’ ability to summarize and explain. One participant, P3, touched on this concept, as well as ChatGPT’s hallucinations (discussed in Section 3.1.2): “I’m a little bit wary of asking it for expert questions. […] [I] asked for it to review a paper for me, and it ended up not actually reading the paper, [and] making generalizations […] I even asked for it to make a work cited, and it gave me cited resources that didn’t actually exist.
3.2. RQ2: How have LLMs influenced software development processes?
Figure 2. Findings related to RQ2, excluding those themes found related to the implementation phase, which can be found in Figure 3.
To answer RQ2, we loosely organize our findings based on the software development life cycle (Services, nd) and the agile development methodology (Fowler et al., 2001). The key themes that emerged in understanding how LLMs affect developers’ software development processes are as follows:
3.2.1. Requirements and Planning
We use a common definition of requirements, defining it as a software capability that must be met by a system or system component in order to satisfy a specification (Leffingwell and Widrig, 2000). Similarly, we define planning as the process of collecting requirements from stakeholders, scheduling, and resource estimation/allocation (Services, nd).
@@ 3.2.1. Requirements and Planning
  • Prototyping: Prototyping was another task that two participants found LLMs to be proficient in. As P12 shared, “[ChatGPT] can […] kind of prototype out what I want to build faster than I could probably do it myself. And then on, when I think it’s looking right or something, I’ll go actually try to implement it.
  • Refining Requirements: LLMs’ assistance in refining requirements, especially for independent contractors, appeared in two interviews. For these participants participants, LLMs did seem to be particularly helpful in this area. P9 shared, “Say, if I got a new contract job that I’m scoping to try to help my clients to try to refine the technical requirements, and especially so on the domains I’m not familiar with. So, most recently, one of the contracts I got was to try to create a music visualizer […] I have never previously dealt with creating apps with specifically musicians before. So while I was talking to them getting requirements at the same time, I would have the ChatGPT window on the side.” P10 shared that ChatGPT helped him research and cut down discovery time when interacting with stakeholders, and that this was important for him throughout the software engineering process: “Because I’m a kind of a one-man shop currently, I really need someone or something to help me to kind of prop up all those processes.
@@
  • Inability to Replace Human Involvement: Twelve participants stated that LLMs did not have an impact on their requirements gathering. Participants mentioned that they received requirements and plans from their supervisors or superiors, who work directly with client needs, and these can’t be generated by LLMs. As P2 shared, “[Requirements gathering] hasn’t changed because most of the work that I’m doing is with client developers and product managers from the client side. So we need to talk to them for the requirements we need to specify — like, we need to gather all the requirements from them. And because our projects are very client-specific, I cannot just go on the Internet to see what they would expect. We need to talk to them to figure out the requirements.
  • Limited Requirements Detailing: One participant, P3, shared how ChatGPT cannot help to generate detailed requirements: “I feel as though the requirements it gives me are very common sense, if that makes sense. But it doesn’t really get into the usability aspects of it, or it doesn’t get into like some of the more fine-tuned [requirements], like how should something be done.
  • Struggles with multi-criteria decision-making: We found in one of our interviews that multi-criteria decision-making can be challenging for LLMs. P16 shared, “So I was looking at flight from [point A to point B]. The flight price with [one airline] went up by a lot of money, and I did not want to pay that much, so I was willing to take different flights, like, you know, fly to [point C], and then [point B], see whatever is cheaper […] and then I said, ‘Okay, so ChatGPT is not good with math or logistics,’ and I did have to do it manually.
3.2.2. Design and Ideation
We consider design to be the process of finding solutions and creating a more detailed technical plan as a result of the requirements-finding process (Services, nd), where ideation is a critical component in designing. Decomposition is also part of the design process, and Chattopadhyay et al. have noted how developers employ particular strategies in order to decompose their tasks into smaller units (Chattopadhyay et al., 2019).
  • Increasing Problem Decomposition: Eight participants acknowledged that LLMs encouraged or necessitated problem decomposition into smaller, more manageable components. This is because LLMs can only take in a limited amount of information at a given time. Consequently, participants needed to either decompose their problems for the LLMs or provide their already decomposed problems. P4 shared that for him, ChatGPT “might force me to break [problems] down […] I guess it might be a good forcing function to make me break them down in the subtasks and subrequirements […] If I can’t explain it to ChatGPT, then it might be an indicator that I don’t yet understand it.
  • Utilizing for Ideation: Six participants acknowledged using LLMs for ideation, leveraging their extensive knowledge bases and generative capabilities. Participants highlighted LLMs’ abilities to generate diverse ideas, albeit with varying levels of accuracy. Because of LLMs’ broad bases of knowledge and generative capabilities (further discussed in Section 3.1.2), they could be relied on as a source of ideas. P12 shared that for them, “Probably the ideation stage is where [Bard] shines the most, because you can’t trust it to be accurate or exactly what you want, but you can trust it to generate some cool content for you, some cool code ideas, some cool writing ideas.
@@ 3.2.2. Design and Ideation
  • Inapplicable for Decomposition: Seven participants shared that LLMs have not affected their problem decomposition, viewing that as inherently reliant on individual reasoning than LLMs can do. P13, for instance, stated, “ I don’t think [it has impacted decomposition]. I like planning them myself and then going to ChatGPT for more fine-grained planning about how I should actually implement the code.” P16 shared that she intentionally did not use ChatGPT for “super big things that need to be broken up,” instead only using it for smaller tasks that she believed would be more likely to be answerable. Other participants alluded to only using LLMs for simple tasks and thus avoided struggling with the decomposition process.
  • Stubborn Responses: Six participants noted that LLMs were stubborn, which could limit their usefulness while designing. P1 noted that “These language models…have a tendency to be stubborn,” and P12 stated that ChatGPT tends to “stay with the previous output a lot.” Another participant, P11, remarked on ChatGPT’s struggles to adapt to new input and modify its outputs: “[ChatGPT] doesn’t do a good job of changing itself.” However, P3, who did not believe ChatGPT was stubborn, shared that his opinion was such because he recognized its inherent limitations with regard to complex tasks: “I’ve never really thought [ChatGPT is] stubborn […] I guess I’ve kind of learned through searching online […] to get it to do like what you want. And then realizing it doesn’t do it, […] it’s not that it doesn’t want to, it just can’t.
3.2.3. Implementation
Figure 3. Findings related to the implementation phase of software development.
The implementation process with LLMs parallels code reuse. Therefore, we adopted Rosson and Caroll’s classification (Rosson and Carroll, 1996), dividing the process into three distinct phases: (1) Finding Context, which entails prompt engineering and the process of locating pertinent responses; (2) Evaluating Context, which centers on evaluating the accuracy and usability of the generated responses; and (3) Integration, which pertains to incorporating the generated text into an individual’s own code or within an existing codebase. We note that the following bullet points’ titles are framed from the perspectives of the developer/prompter, rather than the perspectives/capabilities of the LLMs.
@@ 3.2.3. Implementation
  • Varying Prompt Specificity to Impact Answers. Developers’ diverse preferences for broad prompts, specific prompting, or adjusting query specificity to influence LLMs’ output emerged in fifteen interviews.
@@
    Broad questions are favored when seeking varied answers, avoiding assumptions, or seeking low-effort responses. Participants used broad questions and vague prompts when they were looking for varied answers. P15 shared that while he originally had used specific prompts, he found that using broader prompts resulted in improved outcomes: “[I]t started to give me actually better results, because it knew things that I didn’t know […] When I have a more broad [prompt], then it’s able to kind of formulate its own deduction and get to the problem that it thinks that it’s solving.” P4 similarly shared that, while he had originally used more specific prompts, he found vague prompts to ultimately be more productive: “I’ve been moving to shorter and shorter, more vague prompts. So I used to be, you know, trying to do the few-shot learning approach. […] I never do that anymore. It is such a waste of time.
    Specific prompts are preferred when seeking particular answers, especially in software engineering, or when participants have a clear understanding of what they are asking for. As P2 stated, “I think it should be very specific […] If you are going into other [non-software] domains, where there is no one correct answer, you can go for broad questions. But for software development, I think you need to be very specific.” P16 mentioned, “I think it depends on how well you understand the question you’re asking, because I’ve asked ChatGPT very specific questions, but those were specific and to the point.
    Four participants acknowledged that there is not a single correct strategy for choosing the level of specificity. They emphasized the importance of adaptability and experimentation, suggesting that users can benefit from trying both vague and specific prompts. P7 shared two use cases demonstrating the value of both prompting strategies: “Sometimes I want the question be really generic. For example, I want to design a system. […] I want to look into all those possibilities. In that case, I want to give some general question like, ‘Could you give me some high-level architecture for this system?’ But sometimes I want some answer [to be] really specific. For example, when I read a book, there’s a sentence I don’t really understand why the logic [is] this way. I don’t want [ChatGPT] to give me an approximate answer. I want [a definite] answer.” Being able to determine when to be more or less specific is seen as a valuable skill when using LLMs.
@@
  • Enhancing Accuracy and Clarity with Follow-Up Queries: Fifteen participants shared their strategies of prompting iteratively until getting the desired response. Seven participants shared that follow-up queries improve accuracy, and two shared that they improve clarity. As P16 stated, “I think the follow-up questions allow the users to get clarity.” She elaborated on the importance of follow-ups: “The follow-up questions actually might be more important than [the initial question], more probably equally important. But I get more of my answers from the follow-up question.
@@
  • Unique Prompting Strategies: Prompt engineering as a discipline has become increasingly popular, with courses and guides from all corners of the Internet, including from OpenAI itself (OpenAI, nd). Fourteen participants acknowledged the crucial role of prompt engineering in eliciting accurate and relevant responses from LLMs. P9 shared that he could find a good answer “[a]s long as the prompt is crafted carefully, and the problem is common enough to find a solution.” This ties into results discussed in Section 3.1.2, showing that ChatGPT performs best when it encounters less novel problems.
    Participants demonstrated a willingness to try different prompting strategies, with inspiration from online resources and previous interactions with LLMs. P10 shared his unique strategy of adding and removing context via particular prompting by using ’+’ and ’-’ within his prompts, as well as labeling points (e.g., A, B, C) to refer back to them later.
@@
  • Generalizing Prompts and Code for Security: Thirteen participants employed generalized prompts or used case-specific information rather than project-specific details when interacting with LLMs. P2 shared that, “Because I cannot put all the information into ChatGPT, because it’s very client-specific information, it’s confidential information […] I just give it use cases that are similar.” By focusing on broader contexts or specific use cases, participants aimed to minimize the risk of exposing confidential project information while still obtaining relevant responses from LLMs. This approach allowed participants to strike a balance between leveraging LLMs’ capabilities and safeguarding sensitive data. P5 felt more secure in using this approach: “I think that what makes me comfortable about [ChatGPT’s security] is that […] the prompts that I ask are sort of general or generalized […] and really can’t be tied to any particular person’s identity, or any sensitive piece of information.
  • Providing Examples: Nine participants practiced few-shot prompting with ChatGPT. Few-shot prompting, which has been shown in prior research to have mixed success, entails providing examples to an LLM tool in order to receive a particular output (Reynolds and McDonell, 2021). For example, P11 shared that, “If you’re able to give it an example of what you want, then it can do a better job.
  • Applying Contextual Input for Improved Responses: Nine participants highlighted the significance of providing context when interacting with LLMs, i.e., more context generally leads to better responses from LLMs. Examples include P1’s observation that “The more context, the better,” and P15’s explanation that “The more information that you give it, the better [of] a response it’s going to be.” Two participants discussed how ChatGPT’s ability to retain context distinguishes it from traditional search engines. They noted that this feature simplifies the research process and allows for more effective interactions, as ChatGPT can keep track of previous questions and responses, enabling a more seamless exchange of information. As P14 stated, “It takes a mindset from the Googling part, and the biggest mindset shift is that it keeps context that I can build up on the question.
@@
  • Starting a New Thread for Fresh Answers: Nine participants stated their preference for beginning a new thread when they were dissatisfied with the answers received in a previous interaction with ChatGPT or Gemini. This approach allowed them to refresh the context and seek new responses without the constraints or biases of previous chats. P15 detailed his experience with ChatGPT: [I] have it continue refining what I want to do. Sometimes, it’ll get to the point where it kind of [stops] working out. Maybe the context gets a little skewed the further you get down in the chat. So then I’ll just take whatever it had there and then my problem, and then start a new chat. So, recreate the context. And this is where I’m at and trying to do this, and then I can kind of start over and and refresh where it’s at.” Some participants mentioned that when they chose to open a new one, they often improved the specificity or clarity of their prompts. P16 shared that when she continuously gets an incorrect answer, “I open a new thread because sometimes I’m like, I learned my lesson; I learned. I know what [ChatGPT is] going to say now. So, let me open a new thread and start all over. And I’m going to be more specific, and maybe we can solve this together.
@@
  • Struggles with Integration and Adjustments: Five participants observed that LLMs faced challenges when integrating context and making adjustments or changes. Instances were shared where LLMs struggled to modify specific components of their outputs without affecting the entirety of the generated code. P1 shared an incident in which ChatGPT struggled to change minor components in its output: “It wouldn’t change just that piece of the code. It would change all of the code.” Participants also expressed concerns about LLMs’ abilities to retain the original context over multiple prompts or interactions. One participant, P8, shared that in software engineering (not other disciplines), “[W]hen I’m asking more questions, and I give [ChatGPT] more context, I confuse it more.” He described situations where ChatGPT would forget previous information or modifications, leading to the need to start fresh threads to maintain clarity and accuracy. Similarly, P8 noted that higher specificity in prompts could sometimes result in ChatGPT making assumptions or becoming overwhelmed with additional information. Additionally, P3 shared that ChatGPT often struggled to integrate non-functional requirements: “[ChatGPT] usually kind of messes up on those, which is why I avoid [including] non-functional aspects [when prompting].
  • Limitations of Context Window: Five participants highlighted the limitations of LLMs’ context windows. They noted instances where LLMs ran out of space to generate code or lost context when prompted multiple times. P15 echoed this sentiment of ChatGPT forgetting its original context and how this factors into him opening new threads: “[A]t some point further down the line, it’s going to forget certain things that were really, really important to consider. And so sometimes I’ll say, ‘This is really important to remember.’ […] It’s not always perfect, but that’s kind of why I take the context of everything I learned now up to this point, plus my original message, and then readjust and do a new chat, just so [ChatGPT] can kind of get a fresh start at looking at it.” This limitation could lead to frustration and decreased effectiveness in generating accurate responses, as P12 stated: “I’m probably better off doing it myself. At that point, it’s probably going to be a waste of time to continually trying to re-prompt it and re-prompt it.
  • Necessity of Follow-up Queries: Two participants shared that they often relied on follow-ups. This was in part because LLMs did not always get things right the first time. P9 stated, “Very rarely, […] I can just get the code that I want in one try, even though there are specific prompt engineering templates that I follow.” As mentioned, participants often use broad prompts and then provide specific follow-up questions, demonstrating the iterative process of prompting. One participant, P14, shared that “If it’s an easy function, [the task requires] one [follow-up]. And if it’s a very difficult problem, sometimes it takes like 50.
  • Challenges of Providing Context: While acknowledging the benefits of context, one participant noted the challenges of providing sufficient context, especially when they are unsure of what specific examples to provide. P11 mentioned, “As far as context, [ChatGPT] benefits a lot from examples […] If you can give it an example of what you want, then it can do a better job — but often you don’t know what the example would be, because you’re asking it to do something you don’t know how to do.
@@
  • Verifying through Reading Code: Ten participants shared that they rely on reading their code as a primary method of review. This involves going through the code manually, line by line, to understand its logic, structure, and implementation details. Reading allows developers to identify potential issues, errors, or areas for improvement within their codebase. P9 noted that reading could be a sufficient check for a simple piece of code: “For me, as a professional developer, it is still my job to validate [that] this piece of code is going to work, and it’s going to integrate well into my existing codebase. So to do that, if it’s something that’s simple enough, I can just mentally do a quick check — me reading it line by line — and do a quick check mentally, I can do that.
@@
  • Verifying through Output: Similar to testing, ten participants shared that engage in validating their code by checking the output after reading through the code. This involves examining the results of the code execution, which can provide insights into its correctness and functionality. Methods for output verification include checking console logs, manual testing, and comparing the generated output to expectations. As P8 shared, “If that output would be contradicted with what I expect, I discard it.

### v2

@@ 4.1.3 Supporting Developers’ Personal Growth
  • Enhancing Reassurance, Confidence, and Independence: This theme emerged in six interviews. P2, a newer professional, shared his experience of comparing his answers to ChatGPT’s for corroboration: “It has made the [software engineering] process easier definitely. And the best part is I kind of get the reassurance of a lot of things — that if I’m doing it this way, I can ask it, and I can just check if ChatGPT is giving me a similar answer […], then I’m not doing anything which is majorly wrong.
  • Improving Access to Information: A positive view of LLMs’ ability to ease access to information emerged in three interviews, with participants believing ChatGPT would democratize information access worldwide.
  • Enhancing Job Satisfaction: One participant highlighted ChatGPT’s role in his increased job satisfaction by helping him with tasks he struggled with, allowing him to focus on what he cared about.
@@
  • Inability to Replace Human Decisions: Six participants noted LLMs’ inability to replace human interactions and decisions. For instance, P12 emphasized that software engineering involves more than just coding: “Despite what [LLMs] can do, at the end of the day, it can’t replace — like, what’s difficult about being a software engineer isn’t coding particularly. […] I think what’s difficult about being a software engineer, at least with what I do, is the communication that happens between teams, between coworkers, the email threads, and the chat threads that exist.
@@
  • Concerns about Dependency: Two participants mentioned their preference for solving problems on their own before asking LLMs to avoid over-reliance.
  • Slower Implementation than Humans: One interview noted differences between the time it took humans and LLMs to create code in cases where ChatGPT fails to understand or provide accurate solutions.
4.1.4 Assisting Developers’ Non-Technical Tasks
  • Consulting and Decision-Making: Eleven participants discussed the use of LLMs for consultation or direction. For example, P8 mentioned ChatGPT’s impact on resolving team debates: “[A]lways in technical teams, there exist debates on choosing options, options A and B – both are correct, but which one is better? In this way, we have a judge; we have someone that tells us which approach is better.
  • Summarizing Text and Documentation: Nine participants mentioned the use of LLMs for summarizing different textual data, such as articles, papers, and documentation.
  • Supporting Internal Communication: The use of LLMs to facilitate internal communication by composing documents, aiding in presentations, and assisting in explaining complex tasks to team members emerged in five interviews.
@@ 4.1.4 Assisting Developers’ Non-Technical Tasks
  • Limited Summarizing and Explanation Capabilities: Two participants expressed dissatisfaction with LLMs’ ability to summarize and explain, with one noting ChatGPT’s tendency to make generalizations without actually reading provided materials.
4.2 RQ2: How have LLMs influenced software development processes?
To answer RQ2, we loosely organize our findings based on the software development life cycle [62] and the agile development methodology [63]. We identified impacts across five key phases: (1) Requirements and Planning, where LLM adoption was minimal; (2) Design and Ideation, where LLMs encouraged problem decomposition; (3) Implementation, where developers employed sophisticated prompt engineering strategies; (4) Testing and Code Review, where LLMs showed promise for unit tests but faced limitations; and (5) Debugging, Refactoring, and Documentation, where LLMs reduced debugging time. We organize our findings to follow this progression through the development lifecycle.
4.2.1 Requirements and Planning
We use a common definition of requirements, defining it as a software capability that must be met by a system or system component in order to satisfy a specification [64]. Similarly, we define planning as the process of collecting requirements from stakeholders, scheduling, and resource estimation/allocation [62].
@@ 4.2.1 Requirements and Planning
  • Prototyping and Refining Requirements: Two participants found LLMs proficient in prototyping, and two found them helpful in refining requirements, especially for independent contractors. P9 shared how ChatGPT helped him research and cut down discovery time when interacting with stakeholders: “Say, if I got a new contract job that I’m scoping to try to help my clients to try to refine the technical requirements, and especially so on the domains I’m not familiar with […] while I was talking to them getting requirements at the same time, I would have the ChatGPT window on the side.
@@
  • Inability to Replace Human Involvement: Twelve participants stated that LLMs did not have an impact on their requirements gathering. Participants mentioned that they received requirements and plans from their supervisors or superiors, who work directly with client needs, and these can’t be generated by LLMs. As P2 shared, “[Requirements gathering] hasn’t changed because most of the work that I’m doing is with client developers and product managers from the client side. So we need to talk to them for the requirements we need to specify […] And because our projects are very client-specific, I cannot just go on the Internet to see what they would expect.
  • Limited Requirements Detailing: One participant shared how ChatGPT cannot help to generate detailed requirements, noting that responses were “very common sense” but didn’t address usability aspects or fine-tuned details. Another participant noted struggles with multi-criteria decision-making.
4.2.2 Design and Ideation
We consider design to be the process of finding solutions and creating a more detailed technical plan as a result of the requirements-finding process [62], where ideation is a critical component in designing. Decomposition is also part of the design process, and Chattopadhyay et al. have noted how developers employ particular strategies in order to decompose their tasks into smaller units [65].
  • Increasing Problem Decomposition: Eight participants acknowledged that LLMs encouraged or necessitated problem decomposition into smaller, more manageable components. This is because LLMs can only take in a limited amount of information at a given time. P4 shared that for him, ChatGPT “might force me to break [problems] down […] I guess it might be a good forcing function to make me break them down in the subtasks and subrequirements […] If I can’t explain it to ChatGPT, then it might be an indicator that I don’t yet understand it.
  • Utilizing for Ideation: Six participants acknowledged using LLMs for ideation, leveraging their extensive knowledge bases and generative capabilities. P12 shared that for them, “Probably the ideation stage is where [Bard] shines the most, because you can’t trust it to be accurate or exactly what you want, but you can trust it to generate some cool content for you, some cool code ideas, some cool writing ideas.
@@ 4.2.2 Design and Ideation
  • Inapplicable for Decomposition: Seven participants shared that LLMs have not affected their problem decomposition, viewing that as inherently reliant on individual reasoning. P13, for instance, stated, “I don’t think [it has impacted decomposition]. I like planning them myself and then going to ChatGPT for more fine-grained planning about how I should actually implement the code.”
  • Stubborn Responses: Six participants noted that LLMs were stubborn, which could limit their usefulness while designing. P11 remarked on ChatGPT’s struggles to adapt to new input: “[ChatGPT] doesn’t do a good job of changing itself.
4.2.3 Implementation
The implementation process with LLMs parallels code reuse. Therefore, we adopted Rosson and Carroll’s classification [66], dividing the process into three distinct phases: (1) Finding Context, which entails prompt engineering and the process of locating pertinent responses; (2) Evaluating Context, which centers on evaluating the accuracy and usability of the generated responses; and (3) Integration, which pertains to incorporating the generated text into an individual’s own code or within an existing codebase. We note that the following bullet points’ titles are framed from the perspectives of the developer/prompter, rather than the perspectives/capabilities of the LLMs.
@@ 4.2.3 Implementation
  • Varying Prompt Specificity to Impact Answers. Developers’ diverse preferences for broad prompts, specific prompting, or adjusting query specificity to influence LLMs’ output emerged in fifteen interviews. Broad questions are favored when seeking varied answers or avoiding assumptions. P15 shared that while he originally had used specific prompts, he found that using broader prompts resulted in improved outcomes: “[I]t started to give me actually better results, because it knew things that I didn’t know […] When I have a more broad [prompt], then it’s able to kind of formulate its own deduction and get to the problem that it thinks that it’s solving.” Specific prompts are preferred when seeking particular answers, especially in software engineering. As P2 stated, “I think it should be very specific […] If you are going into other [non-software] domains, where there is no one correct answer, you can go for broad questions. But for software development, I think you need to be very specific.” Four participants acknowledged that there is not a single correct strategy for choosing the level of specificity, emphasizing the importance of adaptability and experimentation.
@@
  • Enhancing Accuracy and Clarity with Follow-Up Queries: Fifteen participants shared their strategies of prompting iteratively until getting the desired response. Seven participants shared that follow-up queries improve accuracy, and two shared that they improve clarity. As P16 stated, “The follow-up questions actually might be more important than [the initial question], more probably equally important. But I get more of my answers from the follow-up question.
@@
  • Unique Prompting Strategies: Fourteen participants acknowledged the crucial role of prompt engineering in eliciting accurate and relevant responses from LLMs. P9 shared that he could find a good answer “[a]s long as the prompt is crafted carefully, and the problem is common enough to find a solution.” Participants demonstrated a willingness to try different prompting strategies, with inspiration from online resources and previous interactions with LLMs.
@@
  • Generalizing Prompts and Code for Security: Thirteen participants employed generalized prompts or used case-specific information rather than project-specific details when interacting with LLMs. P2 shared that, “Because I cannot put all the information into ChatGPT, because it’s very client-specific information, it’s confidential information […] I just give it use cases that are similar.
  • Providing Examples and Context: Nine participants practiced few-shot prompting with ChatGPT, and nine participants highlighted the significance of providing context when interacting with LLMs. P1 observed that “The more context, the better.” Two participants discussed how ChatGPT’s ability to retain context distinguishes it from traditional search engines.
@@
  • Starting a New Thread for Fresh Answers: Nine participants stated their preference for beginning a new thread when they were dissatisfied with the answers received in a previous interaction. P15 detailed his experience: “[I] have it continue refining what I want to do. Sometimes, it’ll get to the point where it kind of [stops] working out. Maybe the context gets a little skewed the further you get down in the chat. So then I’ll just take whatever it had there and then my problem, and then start a new chat.
These six strategies reveal sophisticated prompt engineering practices among early adopters, demonstrating that effective LLM use requires deliberate experimentation, iterative refinement, and careful context management. Notably, participants did not converge on a single ”best practice” but rather developed flexible, context-dependent approaches that balanced specificity with openness, security with informativeness, and persistence with knowing when to start fresh.
@@
  • Struggles with Integration and Adjustments: Five participants observed that LLMs faced challenges when integrating context and making adjustments or changes. P1 shared an incident in which ChatGPT struggled to change minor components in its output: “It wouldn’t change just that piece of the code. It would change all of the code.” Participants also expressed concerns about LLMs’ abilities to retain the original context over multiple prompts or interactions.
  • Limitations of Context Window: Five participants highlighted the limitations of LLMs’ context windows. They noted instances where LLMs ran out of space to generate code or lost context when prompted multiple times.
  • Necessity of Follow-up Queries and Challenges of Providing Context: Two participants shared that they often relied on follow-ups because LLMs did not always get things right the first time. One participant noted the challenges of providing sufficient context, especially when unsure of what specific examples to provide.
@@
  • Verifying through Reading Code: Ten participants shared that they rely on reading their code as a primary method of review. P9 noted that reading could be a sufficient check for a simple piece of code: “For me, as a professional developer, it is still my job to validate [that] this piece of code is going to work, and it’s going to integrate well into my existing codebase. So to do that, if it’s something that’s simple enough, I can just mentally do a quick check — me reading it line by line.
@@
  • Verifying through Output: Similar to testing, ten participants shared that they engage in validating their code by checking the output after reading through the code. As P8 shared, “If that output would be contradicted with what I expect, I discard it.